# Distillation Robustifies Unlearning

### v2

@@ Acknowledgments
We thank Henrik Marklund for very insightful discussions. We are grateful to Rishub Tamirisa for the guidance in navigating WMDP benchmarking procedures and to Eric Easley for sharing valuable strategies for WMDP dataset cleaning and productive discussions about potential improvements to our method. Vivek Hebbar, Andis Draguns, and Jake Mendel offered helpful comments on the abstract.
We thank Iftekhar Uddin and Laura Vaughan for facilitating access to computational resources and funding support. We thank Lisa Wong for help with the design of Figure 1.
Bruce W. Lee’s work on this project was supported by funding from Open Philanthropy and the Long-Term Future Fund. We also thank MATS for facilitating our collaboration and providing the institutional framework that made this research possible.
@@ Appendix A Experimental Details for Section 3
Relearning  To evaluate the robustness of unlearning methods, we conduct relearning experiments to assess whether unlearned capabilities could be easily recovered. Notably, fine-tuning is often considered a strong form of latent capability elicitation, so it can provide insights into whether knowledge is truly removed from the model or just suppressed during oracle matching. For all relearning experiments, we use the final student model from the oracle matching procedure as the starting point. Training proceeds for 500 steps, with validation performed every five steps to track both retention and recovery of capabilities. Table 4 presents the hyperparameters used for our relearning experiments.
[…]
@@ Appendix B Experimental Details for Section 4
In Section 4, we report suppress and distill experiments. Many experimental details are similar to what was reported in Appendix A.
@@
Table 5: Suppress Hyperparameters
@@ B.2 Distillation
Distillation  In Figure 4, we report relearning trends for the suppressed, distilled, and gold standard. We use the same dataset that was used to pretrain respective models as distillation data. We also ensure that the distilled model undergoes the same number of update steps as its pretrained counterpart. We perform distillation with the hyperparameters in Table 6. Additionally, we report the distillation trend in Figure 9. To find information about Data Filtering (Gold Standard) models, refer to Appendix A.
@@
Teacher Model Suppress Only [^5] This refers to the GradDiff/MaxEnt/RMU unlearned models, which are also marked as ”Suppress Only” in Figure 4. Suppress Only Suppress Only Suppress Only Suppress Only Suppress Only
@@ Appendix C Experimental Details for Section 5
In Section 5, we introduce UNDO as an approach to leverage random noising as an alternative to full random initialization.
@@
Noising Method  The novel contribution introduced in Section 5 is controlled parameter corruption, allowing us to interpolate between the original model and random noise. This provides a more flexible approach compared to binary choices of keeping original parameters or fully reinitializing them. The algorithm applies a weighted combination of the original parameters and scaled random noise based on two hyperparameters: $\alpha$ controls the proportion of corruption applied, while $\beta$ controls the magnitude of the noise. The pseudocode is as follows:
@@ D.1 Language and Arithmetic Unlearning
  • RMU: Representation Manipulation for Unlearning, which modifies internal model representations for selected layers.
@@ D.2 WMDP
We apply our method to the WMDP [40] benchmark on both the Cyber and Bio domain and compare to existing unlearning methods.
We use Gemma-2-2b LABEL:team2024gemma as the base model.
@@ D.2.2 Datasets
WDMP includes a WMDP forget and retain corpora for each domain. However, it is common practice to use other datasets as well. For example the original WMDP paper introduces RMU [40] using Wikitext [49] as the retain dataset.
@@ D.2.3 Baseline Unlearning Methods
Both RMU and MaxEnt involve optimizing a loss that is the sum of a retain component and a forget component. For WMDP, we parameterize this combination with $\alpha$ such that $loss=\alpha*retain\_component+(1-\alpha)*forget_{c}omponennt$ where $10<\alpha<$ 1.
@@
We run hyperparameters sweeps on alpha and learning rate for both methods, then chose three points across the frontier to use in the experiment. We run four seeds of each method and apply adversaries to each to measure robust forget performance. Of the three points we use the one with the highest retain performance as the teacher for the UNDO models. Ultimately, the models in Table 10 are chosen.
@@ D.2.5 Relearning Adversaries
The diversity of adversaries highlights how each existing method may be robust against some adversaries, but does not deeply remove knowledge and therefore is not robust against all adversaries. For example, repnoise was much more susceptible to the shrink-perturb relearning, while being more robust to other adversaries. Similarly, for the TAR method, the authors report robustness across most adversaries, but select adversaries were completely able to recover performance LABEL:tamirisa2024tamper. While we were unable to find hyperparameters that maintained reasonable retention performance in the time we spent, we expect this to hold true for TAR in all of our settings as well.
@@ Appendix E Compute requirements
Appendix F Author Contribution
@@
Bruce W. Lee was a core contributor to our experimental codebase and was responsible for framing and conducting most experiments reported in the paper, except WMDP. This includes Unlearn-and-Distill and UNDO experiments reported in Figures 1, 2, 3, 4, 5, 6, 7. BL implemented baselines and plotted all figures in the paper. BL wrote the majority of the later sections and assisted in revising the earlier sections.

### v3

@@ Acknowledgments
We thank Henrik Marklund for very insightful discussions. We are grateful to Rishub Tamirisa for the guidance in navigating WMDP benchmarking procedures and to Eric Easley for sharing valuable strategies for WMDP dataset cleaning and productive discussions about potential improvements to our method. Vivek Hebbar, Andis Draguns, and Jake Mendel offered helpful comments on the abstract. We thank Iftekhar Uddin and Laura Vaughan for facilitating access to computational resources and funding support. We thank Lisa Wong for help with the design of Figure 1. We thank four anonymous reviewers for their time spent on our manuscript and for providing constructive feedback.
Bruce W. Lee’s work on this project was partially supported by funding from Open Philanthropy. We also thank MATS for facilitating our collaboration and providing the institutional framework that made this research possible.
Appendices
@@ Appendix A Experimental Details for Section 3
Relearning  To evaluate the robustness of unlearning methods, we conduct relearning experiments to assess whether unlearned capabilities could be easily recovered. Notably, finetuning is often considered a strong form of latent capability elicitation, so it can provide insights into whether knowledge is truly removed from the model or just suppressed during oracle matching. For all relearning experiments, we use the final student model from the oracle matching procedure as the starting point. Training proceeds for 500 steps, with validation performed every five steps to track both retention and recovery of capabilities. Table 4 presents the hyperparameters used for our relearning experiments.
We assume a significantly stronger adversary compared to contemporary works in LLM unlearning, and show that our methods introduced in Sections 4 and 5 are effective across weaker and stronger adversaries. Our Language Relearning experiments train $\sim$8M tokens, and Arithmetic Relearning experiments train $\sim$2M tokens. On top of this, we do an adversarial learning rate search for all relearning runs over {1e-3, 7e-4, 4e-4, 1e-4, 7e-5, 5e-5, 1e-5, 7e-6, 4e-6, 1e-6}. In all relearning experiments (Figure 2, Figure 4, Figure 5), we always mark the adversarially optimal performance (the best performance the adversary could obtain by varying the learning rate) at each validation step in thicker lines.
@@ Appendix B Experimental Details for Section 4
In Section 4, we report unlearn and distill experiments. Many experimental details are similar to what was reported in Appendix A.
@@
Table 5: Unlearning Hyperparameters
@@ B.2 Distillation
Distillation  In Figure 4, we report relearning trends for the unlearned, distilled, and gold standard. We use the same dataset that was used to pretrain respective models as distillation data. We also ensure that the distilled model undergoes the same number of update steps as its pretrained counterpart. We perform distillation with the hyperparameters in Table 6. Additionally, we report the distillation trend in Figure 9. To find information about Data Filtering (Gold Standard) models, refer to Appendix A.
@@
Teacher Model Unlearn Only [^5] This refers to the GradDiff/MaxEnt/RMU unlearned models, which are also marked as ”Unlearn Only” in Figure 4. Unlearn Only Unlearn Only Unlearn Only Unlearn Only Unlearn Only
@@ Appendix C Experimental Details for Section 5
In Section˜5, we introduce UNDO as an approach to leverage random noising as an alternative to full random initialization.
@@
Noising Method  The novel element introduced in Section˜5 is controlled parameter corruption, allowing us to interpolate between the original model and random noise. This provides a more flexible approach compared to binary choices of keeping original parameters or fully reinitializing them. The algorithm applies a weighted combination of the original parameters and scaled random noise based on two hyperparameters: $\alpha$ controls the proportion of corruption applied, while $\beta$ controls the magnitude of the noise. The pseudocode is as follows:
@@ D.1 Language and Arithmetic Unlearning
  • RMU: Representation Misdirection for Unlearning, which modifies internal model representations for selected layers.
@@ D.2 WMDP
We apply our method to the WMDP [42] benchmark on both the Cyber and Bio domain and compare to existing unlearning methods. We use Gemma-2-2b [73] as the base model.
@@ D.2.2 Datasets
WMDP includes WMDP forget and retain corpora for each domain. However, it is common practice to use other datasets as well. For example the original WMDP paper introduces RMU [42] using Wikitext [51] as the retain dataset.
@@ D.2.3 Baseline Unlearning Methods
Both RMU and MaxEnt involve optimizing a loss that is the sum of a retain component and a forget component. For WMDP, we parameterize this combination with $\alpha$ such that $loss=\alpha*retain\_component+(1-\alpha)*forget\_component$ where $0<\alpha<1$.
@@
We run hyperparameter sweeps on alpha and learning rate for both methods, then chose three points across the frontier to use in the experiment. We run four seeds of each method and apply adversaries to each to measure elicited forget performance. Of the three points we use the one with the highest retain performance as the teacher for the UNDO models. Ultimately, the models in Table 10 are chosen.
@@ D.2.5 Relearning Adversaries
We test diverse adversaries, as each existing method may be selectively robust against certain attacks. For example, RepNoise was much more susceptible to the shrink-perturb relearning, while being more robust to other adversaries. Similarly, for the TAR method, the authors report robustness across most adversaries, but select adversaries were completely able to recover performance [71]. We were unable to find hyperparameters that maintained reasonable retain performance within our training budget.
@@ Appendix E Compute requirements
Appendix F Post-Review Reflection: An Alternative Setup to Study Unlearn-and-Distill
We present supplementary experiments conducted during the review period that extend our findings to additional benchmarks and attack vectors. The main concerns raised centered on the generalizability of our findings beyond simplified settings and the computational requirements of training models from scratch. We address these specifically in Appendices F.1 and F.2.
We first discuss an alternative experimental setup that enables studying the unlearn-and-distill method without reinitializing models from scratch. In this section, consider a pretrained model $M$ and a dataset $D$ that we can guarantee $M$ has never encountered during pretraining. Here, we assume the not encountering the dataset $D$ can guarantee that the pretrained model $M$ doesn’t learn the capabilities associated with it.
Given the setup above, we can then finetune $M$ on $D$ to create $M^{\prime}$, apply unlearning methods on $M^{\prime}$ to suppress $D$, then distill $M^{\prime}$ back onto the original model $M$. Since $M$ never possessed latent knowledge of $D$, it serves as an effective substitute for a randomly initialized network with respect to this specific knowledge.
Our observation throughout the paper was that once a network has learned certain capabilities during pretraining, reversing this learning requires a significant corruption of the network parameters (also see concurrent literature [85] that discusses the level of corruption needed for robust unlearning to be more likely). Indeed, the reference (unlearned) models in Figure˜2 and Figure˜4 demonstrated rapid relearning, and we broadly refer to the neural network properties that enable this rapid recovery as “latent traces.” However, if we have data $D$ that was definitively absent from the pretraining corpus, then $M$ contains no such latent traces for $D$. This allows us to emulate the unlearn-and-distill setup from Section˜4 without access to the original pretraining data of $M$.
There are challenges to obtaining an ideal $D$. Language models are essentially large token classifiers that develop circuits to use all tokens in their vocabulary, including those that would appear in $D$. Even if the semantic content of $D$ is novel, the model has seen and predicted its constituent tokens. Therefore, it’s likely that keyword or semantic matching cannot guarantee the absence of latent traces.
Fortunately, this challenge is not new, and we have an established approximation. The TOFU benchmark [50] contains 200 fictional author profiles with question-answer pairs, constructed to be absent from existing corpora. Here, the specific associations between authors and their biographical details are novel. Therefore, we conduct supplementary experiments in Appendices F.1 and F.2 with the assumption that TOFU provides sufficient separation to study the distillation phenomenon in a controlled setting where $M$ effectively starts at random initialization with respect to the forget knowledge.
Figure 12: Training dynamics across finetuning, unlearning, and distillation phases on TOFU. Panels (a-c) show Gemma-2-2B, and panels (d-f) show Phi-1.5.
F.1 TOFU Experiments
Following the framework presented above, we conducted experiments using the three-step protocol.
  • First, we finetuned Gemma-2-2B / Phi-1.5 on the full TOFU dataset, achieving high accuracy on the fictional author questions.
  • Second, we applied GradDiff (gradient ascent on the forget set; gradient descent on the retain set), suppressing the model’s ability to answer questions about these authors while maintaining performance on the retain set.
  • Third, we distilled the unlearned model back onto a fresh Gemma-2-2B / Phi-1.5 base model. We use the full TOFU dataset that was used to finetune the base model, including both forget and retain domains.
Figure˜12 shows the training dynamics across all three phases. We conduct the experiment using the open-unlearning framework [20] and track seven commonly used unlearning metrics. During finetuning (a, d), both models learn the fictional author information, with forget metrics increasing steadily while MIA (membership inference attack) [68] scores rise to near-perfect levels. The unlearning phase (b, e) shows suppression of forget knowledge, with Forget QA probability dropping to near zero within 50 steps while retain performance stabilizes. The distillation phase (c, f) demonstrates selective knowledge transfer. The student models maintain forget metrics similar to the base model that hasn’t seen TOFU, despite being trained on outputs from the full dataset, while retain metrics gradually recover to match the teacher’s performance.
Figure 13: Relearning dynamics on TOFU forget set for (a) Gemma-2-2B and (b) Phi-1.5. Models that underwent unlearning followed by distillation show relearning rates closer to control baselines that never saw the forget authors, while unlearned-only models rapidly recover the suppressed knowledge. Individual metrics are shown in light colors with average trends in bold.
To evaluate robustness, we subjected three models to relearning attacks: the unlearned-only model, the unlearned-and-distilled model, and a control baseline (the original base model that never saw TOFU). Figure˜13 shows the relearning dynamics over 100 gradient updates. The unlearned-only models rapidly recover the suppressed knowledge, approaching the performance levels seen after initial finetuning. In contrast, the unlearned-and-distilled models show substantially slower relearning, with trajectories closer to the control baselines that never possessed the forget knowledge.
These results demonstrate that the distillation robustification phenomenon holds even when using existing models as fresh networks, provided the forget knowledge was absent from their original training. The consistency across both Gemma-2-2B and Phi-1.5 architectures suggests this approach generalizes across different model families.
F.2 Robustness Beyond Relearning Attacks
Prior work established that gradient-based finetuning represents the strongest elicitation technique for recovering unlearned capabilities [31, 49]. Nevertheless, we tested robustness against alternative attack vectors during the review period. Zhang et al. [92] demonstrated that quantization can bring back unlearned knowledge by exploiting numerical instabilities in model weights. We evaluated INT4 quantization attacks using the open-unlearning framework on our TOFU-trained models.
Method Precision Forget Q&A Prob Forget Truth Ratio Model Utility
GradDiff FP16 0.000 0.000 0.495
GradDiff INT4 0.008 0.185 0.242
GradDiff + Distilled FP16 0.000 0.001 0.442
GradDiff + Distilled INT4 0.003 0.017 0.391
Table 14: Quantization attack results on TOFU (Gemma-2-2B). Lower forget metrics indicate better unlearning.
Under INT4 quantization, standard unlearning shows substantial degradation of unlearning performance, with forget truth ratio jumping from 0.000 to 0.185. The distilled models demonstrate better robustness, with forget truth ratio reaching only 0.017. The distilled models also maintain better utility under quantization (0.391 vs 0.242).
These quantization results suggest that distillation creates fundamentally different internal representations rather than merely suppressing outputs. The student network, having never possessed the forget knowledge in its parameter space, lacks the latent structures that quantization exploits in the unlearned-only models. This aligns with our broader hypothesis that distillation acts as a capability filter, allowing only the knowledge actively demonstrated by the teacher to transfer to the student.
Appendix G Author Contribution
@@
Bruce W. Lee was a core contributor to our experimental codebase and was responsible for framing and conducting most experiments reported in the paper, except WMDP. This includes Unlearn-and-Distill and UNDO experiments reported in Figures 1, 2, 3, 4, 5, 6, 7. BL implemented baselines and plotted all figures in the paper. BL wrote the majority of the later sections and later revised the earlier sections.